# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

### v1

@@ IV-B Integrated Mapping
10:   $L\leftarrow\textsc{Zero}(|V|)$ $\triangleright$ Reset Locks
11:   if $i_{w}<i^{\max}_{w}$ then
12:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{weakReb}(G,\Pi^{\prime})$
13:     $i_{w}\leftarrow i_{w}+1$
14:   else
15:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{strongReb}(G,\Pi^{\prime})$
16:     $i_{w}\leftarrow 0$
17:  $\Pi^{\prime}\leftarrow\operatorname{move}(\mathcal{M},\Pi^{\prime\prime})$
18:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
19:   if $J(C,D,\Pi^{\prime})<J(C,D,\Pi)$ then $\Pi\leftarrow\Pi^{\prime}$
20:   if $J(C,D,\Pi^{\prime})<\phi J(C,D,\Pi)$ then $i\leftarrow 0$
21:  else
22:   if $\operatorname{maxImb}(G,\Pi^{\prime})<\operatorname{maxImb}(G,\Pi)$ then
23:     $\Pi\leftarrow\Pi^{\prime},\quad i\leftarrow 0$
Overall Refinement Algorithm. The complete refinement procedure is summarized in Algorithm 4. The algorithm iteratively alternates between unconstrained label propagation and rebalancing (depending on the current imbalance) until the maximum number of iterations $i^{\max}$ is reached. A copy of the mapping $\Pi^{\prime}$ is used and the actual mapping $\Pi$ is only updated if an improvement mapping has been found. The parameter $i^{\max}_{w}$ controls how many consecutive iterations of weak rebalancing are applied before a single iteration of strong rebalancing is triggered.
In each iteration, if the current partition $\Pi^{\prime}$ satisfies the imbalance constraint (Line 5), the algorithm performs unconstrained label propagation to reduce the communication cost (Line 6). Otherwise, the algorithm invokes weak or strong rebalancing to improve the balance (Lines 8158-15). After each iteration, the partition $\Pi^{\prime}$ is updated (Line 16), and the best feasible solution encountered so far is saved.
The process terminates when no significant improvement in either objective quality or balance has been achieved within the iteration limits. The hyperparameter $\phi=0.999$ specified the minimum relative improvement of $J(C,D,\Pi)$ that has to be achieved in each iteration. We set $i^{\max}_{w}=2$ by default, but increase it to $10$ in the final refinement stage of the multilevel pipeline to ensure minimal loss while achieving balance. The maximum number of iterations $i^{\max}$ is set to $12+\operatorname{level}$, where $\operatorname{level}$ denotes the current level in the coarsening hierarchy (the original graph at level 0, the coarser representation at higher levels). At coarser levels, graphs are smaller, allowing more iterations, but at finer levels, graphs are larger, so each iteration is more computationally expensive.
Additional Data Structure. Determining the gain of moving a vertex to its neighboring blocks is essential for both label propagation and rebalancing. While this information can be obtained by iterating over the neighbors of a vertex, it can be inefficient as a vertex with 100 neighbors might only be connected to two blocks. We employ an additional data structure [18] that enables rapid access to the neighboring blocks. Conceptually, it is similar to the CSR representation of a graph, however, instead of storing edges, it stores the neighboring blocks and their the aggregated weight of edges. For each vertex $v$, we initially allocate memory for $|\{\Pi(u)\mid u\in N(v)\}|+2$ entries. The additional capacity accounts for potential future vertex moves. Note that this is less memory than Jet, which allocates $\min\{k,|N(v)|\}$ per vertex. Whenever a vertex changes its block assignment, we attempt to update its connectivity within the existing memory. We update the connectivity of $u$ to its neighbors $v$ as follows. For each $(u,v,w)\in\mathbb{E}$, we subtract $w$ from $v$’s connection to $\Pi(u)$. Should the resulting weight drop to 0, then the slot is freed. Next, we add the connectivity to $u$’s new block $\Pi^{\prime}(u)$ in $v$’s connectivity, if $\Pi^{\prime}(u)$ does not already exist. If this is not possible, because all slots are already occupied, we mark in a bit array, that $\Pi^{\prime}(u)$ could not be inserted into $v$’s adjacency. As a last step we add the weights $w$ to the slot $\Pi^{\prime}(u)$. All these steps utilize a parallel_for loop over the edge list $\mathbb{E}$. If any vertex cannot be updated because not enough slots are available, we trigger a full rebuild of the data structure. Any vertex that exceeded its capacity receives $|\{\Pi(u)\mid u\in N(v)\}|+2$ slots, while all other vertices retain their slots.
V Experimental Evaluation
@@
All of our algorithms111github.com/HenningWoydt/GPU-HeiProMap are implemented in C++ 17 and utilize the Kokkos library [14] version 4.7.0 for the GPU code. It compiles the host side code with GCC 11.4 and the GPU code with Nvidia Cuda compiler 12.6. We use KaFFPa [40] from the KaHIP library version 3.19.0, which is also implemented in C++. We compile with the highest optimization (-O3).
@@
System. The experiments are conducted on a machine that has an Intel® Xeon® w5-3435X, that runs on a max frequency of 3.13.1Ghz. The system features 128 GB of main memory. The used GPU is a NVIDIA® GeForce RTX™ 4090 with 24GB of memory. It has 16.384 Cuda cores with a boost frequency of 2.52.5Ghz. The system runs Ubuntu 22.04.4.
Benchmark Instances. To make our work more comparable, we choose the same benchmark instances as in [44, 15, 24, 43, 47]. They can be seen in Table I. The small instances come from the SuiteSparse Matrix Collection [11]. These represent task graphs from real application. Six graphs were taken from Walshaws’s benchmark archive [45], which holds instances for benchmarking graph partitioning algorithms. The graphs af_shell9, thermal2, nlr are from the matrix and numerical section of the 10th DIMACS Implementation Challenge [5]. The graphs deu and eur are large road networks of Germany and Europe [12]. Graphs rgg23 and rgg24 are random geometric graphs with $2^{23}$ and $2^{24}$ vertices respectively. Each vertex represents a random point in the unit square, and if their distance is below than $0.55\sqrt{\ln n/n}$ they are connected via an edge. Graphs del23 and del24 also have $2^{23}$ and $2^{24}$ vertices, with each vertex representing a point in the unit square. The edges are connected such that the graph is Delaunay triangulation.
TABLE I: Benchmark instance properties.
SuiteSparse Other
Graph $|V|$ $|E|$ Graph $|V|$ $|E|$
cop20k_A 99 843 1 262 244 af_shell9 $\approx$504K $\approx$8.5M
2cubes_sphere 101 492 772 886 thermal2 $\approx$1.2M $\approx$3.7M
thermomech_TC 102 158 304 700 nlr $\approx$4.2M $\approx$12.5M
cfd2 123 440 1 482 229 deu $\approx$4.4M $\approx$5.5M
boneS01 127 224 3 293 964 del23 $\approx$8.4M $\approx$25.2M
Dubcova3 146 689 1 744 980 rgg23 $\approx$8.4M $\approx$63.5M
bmwcra_1 148 770 5 247 616 del24 $\approx$16.7M $\approx$50.3M
G2_circuit 150 102 288 286 rgg24 $\approx$16.7M $\approx$132.6M
shipsec5 179 860 4 966 618 eur $\approx$18.0M $\approx$22.2M
@@
Walshaw
598a 110 971 741 934 wave 156 317 1 059 331
fe_ocean 143 437 409 593 m14b 214 765 1 679 018
144 144 649 1 074 393 auto 448 695 3 314 611
(a) Solution quality
(b) Speedup
Algorithm Phase Small Large GPU-HM I/O 24.9% 47.9% Jet 75.1% 52.1% GPU-HM-ultra I/O 4.8% 13.7% Jet 95.2% 86.3% GPU-IM I/O 29.3% 74.9% Heavy-Edge Matching 0.4% 1.1% Edge-Contraction 0.4% 4.7% KaFFPa 64.8% 6.2% Label Prop. + Bal. 5.0% 12.9% Uncoarsening 0.1% 0.2%
(c) Runtime breakdown
Figure 1: Comparison of solution quality (a) and speedup (b) of our proposed algorithms. Table (c) shows the runtime breakdown per phase.
@@
Experimental Setup. As the hierarchy we choose $H=4:8:\{1,\ldots,6\}$ and as the distance we choose $D=1:10:100$. The imbalance parameter is set to 3% ($\epsilon=0.03$). Each algorithm is run five times with different seeds and the running time and communication cost $J(C,D,\Pi)$ are averaged across the seeds.
@@
Performance Profiles. Performance Profiles [13] enable us to analyze the quality of multiple algorithms. Let $\mathcal{A}$ be the set of all algorithms and let $\mathcal{I}$ be the set of all instances. The quality of algorithm $A\in\mathcal{A}$ on instance $I\in\mathcal{I}$ is $q_{A}(I)$ and $\text{Best}(I)=\min_{A\in\mathcal{A}}q_{A}(I)$ denotes the best quality across all algorithms on instance $I$. For each algorithm $A$, a performance plot shows the fraction of instances, the $y$-axis, for which $q_{A}(I)\leq\tau\text{Best}(I)$, where $\tau\geq 1$ is on the $x$-axis. For $\tau=1$ the plot shows the fraction of instances that each algorithm solved with the best solution. Algorithms that perform well have a high fraction of instances solved with minimal $\tau$ values, they are towards the top-left corner. Sub-optimal algorithms will have small fractions at minimal $\tau$ values, and only achieve higher fractions with greater $\tau$ values.
V-A Results
@@
We will first analyze the results of our GPU-based hierarchical multisection (GPU-HM) and integrated mapping (GPU-IM) approach. Note that Jet [18], which we utilize in GPU-HM, has an ultra configuration which executes 18 refinement iterations instead of only one, to improve solution quality. We abbreviated our hierarchical multisection approach that uses Jet-ultra as GPU-HM-ultra. In the end, we will compare our methods against the established algorithms SharedMap [44] and IntegratedMapping [15], as they are CPU-based equivalent algorithms to hierarchical multisection and integrated mapping.
Comparison. Figure 1 shows the performance profile (a) and the speedup (b) of our algorithms across all instances. GPU-HM-ultra serves as a baseline to measure the speedup as it achieves the lowest communication cost. The default GPU-HM approach reaches a maximum speedup of $6.4$ with a geometric mean speedup of $4.4$. However, the speedup comes with an increase in communication cost. While the ultra variant finds the best solution on 94.794.7% of instances, GPU-HM does so on only 5.35.3%. On average, the communication cost of GPU-HM-ultra is just 0.10.1% above the best solution, whereas GPU-HM is 9.39.3% worse. GPU-IM is the fastest of our algorithms with a maximum speedup of $16$ and a geometric mean speedup of $5.6$. However, it is also the worst in terms of communication cost, as it did not achieve the lowest communication cost on any instance and has on average 17.417.4% higher communication cost than the best solution.
Table 1c reports the percentage of runtime each algorithm spends in its major phases. Results are grouped by graph size: small graphs with fewer than one million vertices and large graphs with more than one million vertices. For the hierarchical multisection approach, most runtime is spent in the partitioning, which gets executed by Jet [18]. Only for large graphs does I/O become significant, accounting for 47.947.9% of the runtime in the case of GPU-HM. The I/O phase includes reading the graph from disk, constructing its data structures, transferring it to the GPU and writing the final partition to disk. Among these steps, loading and constructing the graph dominate the I/O runtime. Overall, reducing the runtime of Jet or employing a faster GPU-based partitioner appears to be the most promising direction for improving performance further. The runtime distribution for GPU-IM follows a similar pattern. The greatest runtime percentages are in the phases that do not utilize the GPU, namely I/O and utilizing KaFFPa [40] for the initial partitioning. For smaller graphs, one could coarsen the graph further to partition a smaller graph, which would reduce the required time. However, this often leads to imbalanced solutions and worse communication cost. As later results will show, GPU-IM already lags behind other CPU-based algorithms in terms of solution quality. We therefore utilize GPU-IM as a very fast but low-accuracy algorithm. For larger graphs, the integrated mapping approach spends most of its runtime in the I/O phase.
@@
Figure 2: Comparison of speedup (left) and performance profile of achieved communication costs (right) of our proposed algorithms GPU-HM-ultra and GPU-IM to CPU-based algorithms SharedMap and IntegratedMapping.
@@
Comparison to CPU-based Algorithms. We compare our algorithms to the CPU-based solvers SharedMap [44] and IntegratedMapping [15], since they performed well in the experimental evaluation of [44]. SharedMap employs the hierarchical multisection approach, just like GPU-HM. For partitioning, it internally uses KaFFPa [40] in either the Fast or Strong configuration, which trade speed for solution quality. We abbreviate the variants by SharedMap-F and SharedMap-S. IntegratedMapping, in contrast, minimizes $J(C,D,\Pi)$ via the multilevel scheme, making it conceptually similar to GPU-IM. It also provides a Fast and Strong configuration, which we denote by IntMap-F and IntMap-S. For comparison, we focus on GPU-HM-ultra, which achieves the highest solution quality, and GPU-IM, which provides the lowest runtime. GPU-HM is omitted for better visual clarity, as its performance lies between the two.
@@
Figure 2 shows the speedup over SharedMap-S (left) and the performance profile (right). Among all algorithms, SharedMap-S achieves the highest solution quality, obtaining the best solution on 125 instances (83.383.3%). However, this comes at the cost of runtime, as SharedMap-S is also by far the slowest algorithm. GPU-HM-ultra finds the best solution on the remaining 25 instances (16.716.7%). All other algorithms are not able to determine a best solution on any instance. Ranked by average additional cost over the best solution, the algorithms are ordered as follows: SharedMap-S (0.020.02%), GPU-HM-ultra (11.511.5%), IntMap-S (14.514.5%), IntMap-F (21.121.1%), GPU-IM (30.730.7%) and SharedMap-F (31.731.7%). GPU-IM and SharedMap-F are at most $2.42\times$ and $2.66\times$ worse than the best-known solution, respectively.
@@
In terms of runtime, GPU-IM is the fastest algorithm overall, achieving a maximum speedup of $598.5$ and a geometric mean speedup of $77.7$ compared to SharedMap-S. While GPU-HM-ultra achieves a maximum speedup of $377$, its geometric mean speedup is only $13.8$. On many instances, it is slower than IntMap-F/S and SharedMap-F, largely due to the partitioning phase being the bottleneck (see Table 1c). The remaining mean speedups are $28.4$ for SharedMap-F, $23.8$ for IntMap-F, and $10.4$ for IntMap-S.
Our algorithms have two clear equal CPU-based counterparts. GPU-HM-ultra slightly outperforms IntMap-S in terms of solution quality. Its geometric mean speedup of $13.8$ also exceeds the $10.4$ of IntMap-S, although it is not faster across all instances. Specifically, IntMap-S is faster on 7171% of small instances (speedup of $1.4$), whereas GPU-HM-ultra dominates on 9696% of the large instances (speedup of $5.6$). Similarly, GPU-IM is comparable to SharedMap-F in terms of solution quality. On small graphs, it achieves on average 2.62.6% lower communication costs, while on larger graphs it performs about 2.32.3% worse. The geometric mean speedup of GPU-IM over SharedMap-F is $1.4$ and $10.6$ for small and large graphs respectively.
@@
In summary, GPU-HM-ultra outperforms the CPU-based algorithm IntMap-S in both solution quality and running time. However, it cannot reach the state-of-the-art solution quality of SharedMap-S. Nevertheless, GPU-HM-ultra is on average an order of magnitude faster. GPU-IM can be almost up to $600$ times faster than SharedMap-S, but its solution quality is on average 30.730.7% higher than the best-known solution.

### v2

@@ 4.2 Integrated Mapping
10:   if $i_{w}<2$ then
11:     $\mathcal{M},\Pi^{\prime\prime}\leftarrow\operatorname{weakReb}(G,\Pi^{\prime})$
12:     $i_{w}\leftarrow i_{w}+1$
13:   else
14:     $\mathcal{M},\Pi^{\prime\prime}\leftarrow\operatorname{strongReb}(G,\Pi^{\prime})$
15:     $i_{w}\leftarrow 0$
16:  $\Pi^{\prime}\leftarrow\operatorname{Move}(\mathcal{M},\Pi^{\prime\prime})$
17:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
18:   if $J(\mathcal{C},\mathcal{D},\Pi^{\prime})<J(\mathcal{C},\mathcal{D},\Pi)$ then $\Pi\leftarrow\Pi^{\prime}$
19:   if $J(\mathcal{C},\mathcal{D},\Pi^{\prime})<\phi J(\mathcal{C},\mathcal{D},\Pi)$ then $i\leftarrow 0$
20:  else
21:   if $\operatorname{maxImb}(G,\Pi^{\prime})<\operatorname{maxImb}(G,\Pi)$ then
22:     $\Pi\leftarrow\Pi^{\prime},\quad i\leftarrow 0$
Overall Refinement Algorithm. The complete refinement procedure is summarized in Algorithm 6. The algorithm iteratively alternates between unconstrained label propagation and rebalancing (depending on the current imbalance) until at least 12 iterations have been executed. This guarantees that at least three full cycles of label propagation, then two weak rebalancing steps and at last one strong rebalancing can be executed. A copy $\Pi^{\prime}$ is used and the actual mapping $\Pi$ is only updated if an improved mapping has been found. In each iteration, if the current partition $\Pi^{\prime}$ satisfies the imbalance constraint (Line 5), the algorithm performs unconstrained label propagation to reduce the communication cost (Line 6). Otherwise, the algorithm invokes weak or strong rebalancing to improve the balance (Lines 8 – 15). After each iteration, the partition $\Pi^{\prime}$ is updated (Line 16), and the best feasible solution encountered so far is saved. The process terminates when no significant improvement in either objective quality or balance has been achieved within the iteration limits. The hyperparameter $\phi=0.999$ specifies the minimum relative improvement of $J(\mathcal{C},\mathcal{D},\Pi)$ that has to be achieved in each iteration. The constants 12, 2 and $\phi$ are transferred from Jet.
During the algorithm, an additional structure (not shown in the pseudocode) stores for each vertex $v$ all neighboring blocks and the sum of edge weights to those blocks. This structure is used to efficiently compute gains during label propagation and rebalancing. For each vertex it holds a hash array of size $\min(|N(v)|,k)$ to store all possible connected blocks. In contrast to Jet, the structure is initialized by using the extended CSR format. By using an edge-parallel loop each thread cycles through the vertex hash array and uses CAS operations to find a valid entry. When moving the vertices (Line 15) the structure must be updated. This is done by either completely resetting the affected vertices’ hash arrays and filling them from scratch, or by first removing old connections and then adding new connections.
5 Experimental Evaluation
@@
All of our algorithms222 / GPU-HeiPa are implemented in C++ 20 and utilize the Kokkos library [15] version 5.0.0 for the GPU code. It compiles the host-side code with GCC 11.4.0 and the GPU code with NVIDIA Cuda compiler 12.6.77. We compile with the highest optimization (-O3).
@@
System. The experiments are conducted on a machine that has an Intel® Xeon® w5-3435X that runs at a max frequency of $3.1$ GHz. The system features 128 GB of main memory. The GPU is an NVIDIA® GeForce RTX™ 4090 with 24 GB of memory. It has 16,384 Cuda cores with a boost frequency of $2.5$ GHz. The system runs Ubuntu 22.04.4.
Table 1: Benchmark instances with their corresponding numbers of vertices and edges.
SuiteSparse Other Walshaw
Graph $|V|$ $|E|$ Graph $|V|$ $|E|$ Graph $|V|$ $|E|$
cop20k_A 99 843 $\approx$1.3M afshell9 504 855 $\approx$8.5M 598a 110 971 741 934
2cubes_sphere 101 492 772 886 thermal2 $\approx$1.2M $\approx$3.7M fe_ocean 143 437 409 593
thermomech_TC 102 158 304 700 nlr $\approx$4.2M $\approx$12.5M 144 144 649 $\approx$1.1M
cfd2 123 440 $\approx$1.5M deu $\approx$4.4M $\approx$5.5M wave 156 317 $\approx$1.1M
boneS01 127 224 $\approx$3.3M del23 $\approx$8.4M $\approx$25.2M m14b 214 765 $\approx$1.7M
Dubcova3 146 689 $\approx$1.7M rgg23 $\approx$8.4M $\approx$63.5M auto 448 695 $\approx$3.3M
bmwcra_1 148 770 $\approx$5.2M del24 $\approx$16.8M $\approx$50.3M
G2_circuit 150 102 288 286 rgg24 $\approx$16.8M $\approx$132.6M
shipsec5 179 860 $\approx$5.0M europe_osm $\approx$50.9M $\approx$54.1M
@@
Benchmark Instances. To make our work more comparable, we choose the same benchmark instances as in [45, 16, 25, 44, 48]. They can be seen in Table 1. The small instances come from the SuiteSparse Matrix Collection [12]. These represent task graphs from real applications. Six graphs were taken from Walshaws’s benchmark archive [46], which holds instances for benchmarking graph partitioning algorithms. The graphs afshell9, thermal2, nlr are from the matrix and numerical section of the 10th DIMACS Implementation Challenge [6]. The graphs deu and europe_osm are large road networks of Germany and Europe [13]. Graphs rgg23 and rgg24 are random geometric graphs with $2^{23}$ and $2^{24}$ vertices respectively. Each vertex represents a random point in the unit square, and if their distance is less than $0.55\sqrt{\ln n/n}$ they are connected via an edge. Graphs del23 and del24 also have $2^{23}$ and $2^{24}$ vertices, with each vertex representing a point in the unit square. The edges are connected such that the graph is a Delaunay triangulation.
@@
Experimental Setup. As the hierarchy, we choose $H=4:8:\{1,\ldots,6\}$ and as the distance, we choose $D=1:10:100$. The imbalance parameter is set to 3% ($\epsilon=0.03$). Each algorithm is run five times with different seeds and the running time and communication cost $J(\mathcal{C},\mathcal{D},\Pi)$ are averaged across the seeds. We do not include the time to initially read the graph from disk and constructing it on the CPU, and also not the time to write the partition to disk. These steps are conceptually the same for all algorithms.
@@
Performance Profiles. Performance Profiles [14] enable us to analyze the quality of multiple algorithms across multiple instances. Let $\mathcal{A}$ be the set of all algorithms and let $\mathcal{I}$ be the set of all instances. The quality of algorithm $A\in\mathcal{A}$ on instance $I\in\mathcal{I}$ is $q_{A}(I)$ and $\text{Best}(I)=\min_{A\in\mathcal{A}}q_{A}(I)$ denotes the best quality across all algorithms on instance $I$. For each algorithm $A$, a performance profile shows the fraction of instances on the $y$-axis, for which $q_{A}(I)\leq\tau\text{Best}(I)$, where $\tau\geq 1$ is on the $x$-axis. For $\tau=1$, the profile shows the fraction of instances that each algorithm solved with the best solution. Algorithms that perform well have a high fraction of instances solved with minimal $\tau$ values, they are towards the top-left corner. Sub-optimal algorithms will have small fractions at minimal $\tau$ values, and only achieve higher fractions with greater $\tau$ values.
5.1 Results
@@
First, we analyze our integrated mapping (GPU-IM) and hierarchical multisection (GPU-HM) algorithms. Jet [19], which is utilized in GPU-HM, also has an ultra configuration, which executes the complete refinement algorithm 18 times. This yields better solution quality at the cost of runtime. We abbreviate our approach that uses the ultra variant as GPU-HM-ultra. In the end, we compare our methods against the established algorithms SharedMap [45] and IntMap [16], which are CPU-based counterparts to hierarchical multisection and integrated mapping.
@@
Figure 1: Performance profile comparing the solution quality (left) and a speedup plot (right) of our proposed algorithms.
5.2 Own Comparison
Figure 1 shows the performance profile and the speedup of our algorithms across all instances. GPU-HM-ultra serves as a baseline to measure the speedup as it achieves the lowest communication cost. The default GPU-HM approach reaches a maximum speedup of $9.1\times$ with a geometric mean speedup of $6.5\times$. However, the speedup comes with an increase in communication cost. While the ultra variant finds the best solution on $95.3\%$ of instances, GPU-HM does so on only $4.7\%$. On average, the communication cost of GPU-HM-ultra is just $0.2\%$ above the best solution, whereas GPU-HM is $5.1\%$ worse. GPU-IM is the fastest of our algorithms with a maximum speedup of $150.1\times$ and a geometric mean speedup of $64.9\times$. However, it did not achieve the lowest communication cost on any instance and has, on average, $17.4\%$ higher communication cost than the best solution.
Table 2 reports the average percentage of runtime GPU-IM spends in its major phases. The hierarchical multisection approaches spend more than $95\%$ of runtime in the partitioning, so we do not analyze it any further. The results in the table are grouped by graph size: small graphs with fewer than one million vertices and large graphs with more than one million vertices. For smaller graphs, the most time consuming phase is the refinement phase where the algorithm almost spends two-thirds of the runtime. Both the coarsening and the initial partitioning phase take about $13\%$ of runtime, and the other phases have a negligible impact on the running time. For larger graphs the distribution is more balanced. Notably, refinement and rebalancing is again the most time consuming phase, with about $45\%$ of total runtime. The initial partitioning now also becomes negligible, however coarsening and contraction take about $11\%$ each. Interestingly all miscellaneous operations, which also include uploading and downloading the initial graph and the resulting partition from the CPU to the GPU and vice versa, is the second greatest runtime consumption. This phase is mostly hardware dependent and cannot be sped up by further algorithm improvements. For the largest and smallest graphs, namely cop20k_A and europe_osm, the absolute running time of each phase is shown in milliseconds for the hierarchy $4:8:6$.
Table 2: Runtime distribution across GPU-IM algorithm phases. The last two columns hold the absolute running times of each phase in milliseconds.
Phase Small Large cop20k_A (ms) europe_osm (ms)
Coarsening 13.02% 11.59% 4.351 41.020
Contraction 3.49% 11.16% 1.010 38.694
Init. Part. 13.85% 4.23% 11.116 7.244
Uncontr. 0.14% 0.24% 0.046 1.523
Refine + Reb. 65.22% 45.53% 24.359 116.598
Misc 4.28% 27.24% 1.193 115.469
Total 100% 100% 42.076 320.550
5.3 CPU-based Comparison
@@
We compare our algorithms to the CPU-based solvers SharedMap [45] and IntMap [16], since they performed well in the experimental evaluation of [45]. SharedMap employs the hierarchical multisection approach, just like GPU-HM. It internally uses KaFFPa [41] for its graph partitioning. IntMap, in contrast, minimizes $J(\mathcal{C},\mathcal{D},\Pi)$ via the multilevel scheme, making it conceptually similar to GPU-IM. Both algorithms have a Fast and Strong configuration which we denote by the suffix -F and -S, respectively. These configurations trade speed for solution quality. For comparison, we focus on GPU-HM-ultra, which achieves the highest solution quality, and GPU-IM, which provides the lowest running time. GPU-HM is omitted for better visual clarity, as its performance lies between the two.
@@
Figure 2 shows the speedup over SharedMap-S and the performance profile. Among all algorithms, SharedMap-S achieves the highest solution quality, obtaining the best solution on 124 instances ($82.7\%$). However, this comes at the cost of runtime, as SharedMap-S is also by far the slowest algorithm. GPU-HM-ultra finds the best solution on the remaining 26 instances ($17.3\%$). All other algorithms are not able to determine a best solution on any instance. Ranked by average additional cost over the best solution, the algorithms are ordered as follows: SharedMap-S ($0.2\%$), GPU-HM-ultra ($12.2\%$), IntMap-S ($14.4\%$), IntMap-F ($20.9\%$), SharedMap-F ($30.8\%$), and GPU-IM ($33.1\%$). SharedMap-F and GPU-IM are at most $1.99\times$ and $3.07\times$ worse than the best-known solution, respectively.
Figure 2: Comparison of the speedup (left) and the performance profile of achieved communication costs (right) of our proposed algorithms GPU-HM-ultra and GPU-IM to CPU-based algorithms SharedMap [45] and IntMap [16].
@@
In terms of runtime, GPU-IM is the best algorithm overall, achieving a maximum speedup of $12376.9\times$ and a geometric mean speedup of $1454.6\times$ compared to SharedMap-S. In fact, it is the fastest algorithm on all 150 instances. While GPU-HM-ultra achieves a maximum speedup of $934.7\times$, its geometric mean speedup is only $22.4\times$. The remaining mean speedups are $42.7\times$ for SharedMap-F, $36.7\times$ for IntMap-F, and $11.7\times$ for IntMap-S.
Our algorithms have two clear CPU-based counterparts. GPU-HM-ultra outperforms IntMap-S in terms of solution quality and speed. It finds better solutions on 117 out of 150 instances ($78.0\%$) and is also faster on 93 instances ($62.0\%$). When split into small and large graphs, with fewer and more than one million vertices respectively, GPU-HM-ultra is only faster on 45 out of 102 small instances, but faster on all 48 large instances with a mean speedup of $12.6\times$. When inspecting the communication cost, GPU-HM-ultra has better solution quality on 95 small instances, while only being better on 22 large instances. Similarly, GPU-IM’s solution quality lies between SharedMap-F and IntMap-F. On smaller graphs GPU-IM finds better solutions than SharedMap-F on 70 instances, but only on 5 instances when compared to IntMap-F. For larger graphs this ratio worsens as GPU-IM only finds better solutions on 14 and 10 instances when compared to SharedMap-F and IntMap-F respectively. However, GPU-IM is much faster than both algorithms with a mean speedup of $55.8\times$ and a max speedup of $221.6\times$ over SharedMap-F. Compared to IntMap-F, the mean speedup is $70.8\times$ and the max speedup is $243.1\times$.
@@
In summary, GPU-HM-ultra outperforms IntMap-S in both solution quality and running time. While being about $20$ times faster than SharedMap-S, it cannot reach its state-of-the-art solution quality. GPU-IM achieves a mean speedup of $1454\times$ compared to SharedMap-S, but its communication cost is on average $33.1\%$ higher. Nonetheless, its quality is comparable to IntMap-F and SharedMap-F while being about $60$ times faster.
5.4 GPU-based Comparison